(9) Let E R and let A E L(R). Define a map f : R -> R by f (x) A,)v. Here (is the Euclidean inner product (a) Prove that
Answer #1

(a) The map f is a polynomial in x_i where x = (x_i)_{n \times 1} . Hence it is a C^1 function.

To find the derivative of f we look at partially differentiating it with respect to the variables. That is

f^{\prime}(x) = \Bigg( \dfrac{\partial f}{ \partial x_1} , \dfrac{\partial f}{ \partial x_2} , \dots , \dfrac{\partial f}{ \partial x_n} \Bigg)

Now  \dfrac{\partial f}{ \partial x_i} = \dfrac{\partial}{ \partial x_i} (x+\langle Ax,x \rangle v)

= \dfrac{\partial}{ \partial x_i} (x) + \dfrac{\partial}{ \partial x_i} (\langle Ax,x \rangle v) = \dfrac{\partial}{ \partial x_i} (x) + \langle\dfrac{\partial}{ \partial x_i} Ax, x \rangle + \langle Ax, \dfrac{\partial}{ \partial x_i} x \rangle

= e_i + \langle \begin{pmatrix} a_{1i} \\ .\\ .\\ .\\ a_{ni} \end{pmatrix} , x \rangle v + \langle Ax, e_i \rangle v = e_i + \sum_{k=1}^n x_k (a_{ki} + a_{ik})) v

Where e_i is a column matrix with 1 in the ith place and 0 else where and A = (a_{ij})

This is the derivative of f at x.

(b) At x=0, f(0) = 0 and the derivative f^{\prime}(0) = \Big( e_1, e_2, \dots , e_n \Big) = I . I is invertible hence by inverse function theorem, there are neighborhoods U and V of the origin such that f : U \rightarrow V is invertible. Therefore a bijection.